# EvolveNav: Self-Improving Embodied Reasoning for LLM-Based Vision-Language Navigation

### v2

@@ 4.3 Ablation Study
Effect of different method components. Table 3 presents the ablation study on Val Unseen set on R2R, where we can find the effectiveness and reasonability of different method components in our EvolveNav. From Table 3, we can observe that through the Stage 1 training of formalized CoT supervised fine-tuning (SFT), the model’s navigational reasoning ability can be significantly enhanced to improve the navigation performance. The introduction of self-enriched CoT labels and the self-reflective auxiliary task can further bring performance gain in both SR and SPL metrics compared to pure CoT SFT. Our full model achieves the best results in SR and SPL among all variants, demonstrating the effectiveness of our two-phase self-improving embodied training framework.
Table 3: Ablation study of method components on Val Unseen set on R2R. We adopt the imitation learning (IL) setting for evaluation.
@@
Baseline 46.90 54.63 43.78
   + Formalized CoT SFT 49.62 59.44 46.26
   + Self-Enriched CoT SFT 50.47 57.48 47.98
   + Self-reflective task 50.51 57.74 47.86
   + Full Model (ours) 51.15 59.18 48.27
Effect of constructed CoT labels. Table 4 compares the navigation performance under different kinds of CoT labels, where we can find the effectiveness of our introduced CoT labels by predicting landmark and direction information in a formalized way. To realize “Free-form CoT”, we introduce the free-form CoT labels collected in NavGPT-2 zhou2024navgpt to train the navigation agent. To realize “Direction & Landmark”, we obtain the best matched landmark in the instruction to each ground-truth observation through the CLIP model radford2021learning . The comparison between “Free-form CoT” and “Direction & Landmark (ours)” shows that our formalized CoT labels can effectively reduce redundant reasoning information to improve the navigational reasoning and decision accuracy. Moreover, the inference time to generate CoT at one timestep of “Free-form CoT” is $\sim$7.8s compared to $\sim$2.5s of “Direction & Landmark (ours)”, demonstrating that our method can significantly promote the reasoning speed ($\sim$×$\times$3 improvement), which is crucial for sequential decision making task like navigation. By comparing “Only Direction” and “Only Landmark” with “Direction & Landmark (ours)”, we can find that both landmark and direction information are important to guide navigation decisions, revealing the reasonability of our constructed CoT labels. The superiority of “Direction & Landmark (ours)” over “Direction & Landmark” demonstrates that our introduced CoT labels, which contain diverse landmarks, can potentially encourage the navigation agent to learn cross-modal alignment knowledge to accurately align the visual observation to the navigation instruction.
Table 4: Ablation study of CoT on Val Unseen set on R2R. denotes using the landmark mentioned in the instruction.
@@ 4.4 Visualization
Fig. 3 gives the action decision visualization comparison between NaviLLM zheng2024towards and our EvolveNav, where we can find that EvolveNav generates reasonable navigational reasoning about landmarks and directions to guide correct action decision making. For example, when $t=0$, from the observations, EvolveNav infers that an observation with door and hallway represents the exit from the bedroom while NaviLLM mistakenly selects an action that remains in the bedroom. Another example is in the hallway ($t=4$ for NaviLLM and $t=1$ for EvolveNav), EvolveNav generates reasoning consistent with the decision of entering the correct doorway on the left. However, NaviLLM chooses the wrong doorway to go back. These results highlight the effectiveness of our approach in improving navigational reasoning for accurate instruction understanding and action prediction.
Figure 3: Action decision visualization of NaviLLM zheng2024towards and our EvolveNav. We only extract two steps and display local candidate space for simplicity. Observations selected by EvolveNav (also are the ground-truth actions) and NaviLLM are annotated by green boxes and red boxes, respectively.
5 Conclusion
In this work, we propose EvolveNav, a novel self-improving embodied reasoning framework to fulfill generalizable and adaptable reasoning for enhancing LLM-based vision-and-language navigation. Through introducing the formalized CoT supervised fine-tuning and self-reflective post-training in the proposed framework, the agent’s navigational reasoning ability can be effectively enhanced while mitigating the overfitting to the training reasoning label distributions simultaneously to improve generalization. Experimental results on two popular benchmarks R2R and CVDN reveal the promising capability of our method in boosting the reasoning ability and decision accuracy for LLM-based navigation agents. We believe that our EvolveNav can provide meaningful references for designing self-improving embodied reasoning paradigms to benefit future LLM-assisted Embodied AI research.
Supplementary Material
Appendix A Overview
In these supplementary materials, we first analyze the broader impacts and limitations of our EvolveNav in Sec. B and Sec. C, respectively. Then, we describe more implementation details in Sec. D. Finally, we present more experimental results, including the landmark extraction visualization, self-enriched CoT label visualization, and loss & performance curves during Self-Reflective Post-Training, in Sec. E.
Appendix B Broader Impact
The scalable deployment of LLM-based navigation robots holds significant potential to enhance operational efficiency in public spaces, reduce human labor costs, and benefit diverse populations including the elderly and individuals with disabilities. Our proposed self-improving embodied reasoning framework can help improve both navigation success rate and decision interpretability of LLM-based navigation agents, therefore contributing to scalable application of LLM-based navigation robots which have high robustness and reliability. However, the inherent hallucination of LLMs may cause incorrect decisions in complex environments, leading to potential risks such as collisions. More efforts devoted to hallucination suppression of LLMs and safety promotion of navigation robots are encouraged for future research in the community.
Appendix C Limitations
(1) In the Stage 1 training of Formalized CoT Supervised Fine-tuning in our EvolveNav, we conduct full-parameter training for the open-source LLM, vicuna-7B [1], like the baseline method [66]. This process costs $\sim$1.5 day training on 8 NVIDIA A100 GPUs with 80G memory, which is relative resource-intensive compared to conventional VLN approaches [10, 9, 17]. In future work, we plan to introduce parameter-efficient training strategies or resort to more lightweight LLMs to reduce the computational resource consumption.
(2) During the formalized CoT label collection for supervised fine-tuning, we employ the open-vocabulary image captioning model [24] to obtain the caption for visual observations, followed by the NLP tool Spacy [20] to extract the landmark from the caption. This process is performed in an offline manner, which might introduce noise for the CoT label construction due to the possible inaccuracy of the image captioning model and the NLP tool. Future work includes utilizing more powerful open-vocabulary models or conducting human checking to promote the accuracy of the CoT labels.
Appendix D Implementation Details
We adopt Vicuna 7B [1] as our large language model (LLM) backbone, following [66]. For R2R [5], we also introduce the auxiliary tasks of trajectory summarization and embodied question answering in [66] during training. The maximum number of landmark extracted for constructing formalized CoT labels is set as 5 to retain useful information and simultaneously promote the reasoning speed. To mitigate the possible landmark missing during landmark detection due to the view deviation for a single candidate observation, we fuse the landmark information of candidate observations with the same direction to construct the CoT labels. We remove some uninformative landmarks such as “floor” and “ceiling” empirically during the landmark extraction. During Stage 1 training, we introduce the CoT supervised finetuning loss $\mathcal{L}_{\mathrm{SFT}}$ under a probability of 0.5 to mitigate the overfitting to the pre-constructed CoT labels. The maximum numbers of training steps for Stage 1 and 2 are set as 60000 and 9000 steps, respectively. Training for Stage 1 with full parameter lasts for $\sim$1.5 days with $\sim$73G GPU memory, and training for Stage 2 with LoRA lasts for $\sim$1 day with $\sim$30G GPU memory. The inference is conducted on 4 V100 GPUs with the evaluation batch size of 2. During inference, we set the LLM temperature to 0 and opt for a greedy strategy to generate action decisions.
Appendix E More Experimental Results
Landmark extraction visualization. We compare different methods for constructing the formalized CoT labels, to verify the reasonability of our landmark extraction strategy by combining image captioning model [24] with the NLP tool Spacy [20]. Fig. 4 presents the landmark extraction visualization comparison of three image captioning models, Tag2Text [24], BLIP-v2 [29], LLaVA 1.6 vicuna 13b [35], and an open-vocabulary object recognition model RAM [64]. Concretely, we provide four candidates in a navigational step for these methods and use Spacy [20] to extract the landmarks in the output caption (except for RAM of directly obtaining tagging). From Fig. 4, we can observe that Tag2Text can capture more informative landmarks while having less redundancy and illusion. For example, for “Cand 2”, Tag2Text correctly detects a kitchen and living room, a dining table, and chair, while BLIP-v2 only detects a large dining table. Although LLaVA 1.6 13b generates abundant captions, it brings noisy landmarks like the outdoor after noun phrases extraction. RAM also generates meaningless and non-existent landmarks, like floor and armchair. These results show that our combination of Tag2text model [24] and NLP tool for landmark extraction can effectively retain informative landmarks while reducing redundancy and illusion for constructing CoT labels, which can enable the agent to better learn cross-modal alignment between observations and instructions.
Figure 5: Self-enriched chain-of-thought (CoT) label visualization. Newly introduced landmarks in the self-enriched CoT label are highlighted in red fonts. GT action denotes the ground-truth action (observation). We omit the direction information in the CoT label.
@@ Appendix E More Experimental Results
Loss & performance variation during Stage 2 training. Fig. 6 shows the loss and performance curves during Self-Reflective Post-Training (Stage 2). In Fig. 6, we can find that our self-reflective loss $\mathcal{L}_{\mathrm{sr}}$, the total training loss $\mathcal{L}_{\mathrm{Stage2}}$, and the navigation error (NE) have similar variation trends. Especially, both two loss curves and the NE curves achieve the lowest value around epoch 9. These results show the effect of our self-reflective auxiliary task during Self-Reflective Post-Training in improving the navigational reasoning and decision accuracy of the LLM-based navigation agent.

### v3

@@ 4.3 Ablation Study
Effect of different method components. Table VI presents the ablation study results on Val Unseen set on R2R, where we can find the effectiveness and reasonability of different method components in our EvolveNav. From Table VI, we can observe that through the Stage 1 training of formalized CoT supervised fine-tuning (SFT) (“1”), the model’s navigational reasoning ability can be significantly enhanced to improve the navigation performance. In Stage 2 training, the introduction of self-enriched CoT labels (“2”) and the self-reflective auxiliary task (“3”) can further bring performance gain respectively in both SR and SPL metrics compared to pure CoT SFT in Stage 1. Our full model (“Full Model”) achieves the best results in SR and SPL compared with “2” and “3”, demonstrating that the combination of our self-enriched CoT label enhancement strategy and self-reflective auxiliary task is non-trivial. Especially, rather than pure self-reflective auxiliary task (“3”) that asks the agent to learn to distinguish fixed correct and wrong reasoning patterns like conventional auxiliary task design, the combination of our self-enriched CoT label enhancement strategy and self-reflective auxiliary task (“Full Model”) can encourage the agent to learn diverse correct reasoning and therefore increase its generalization ability to unseen scenarios.
TABLE VI: Ablation study of method components on Val Unseen set on R2R. We adopt the imitation learning (IL) setting for evaluation. “CoT SFT” represents the Stage 1 training of Formalized CoT Supervised-Finetuning. “Self-Enriched CoT SFT” denote the Self-Enriched CoT Label Enhancement strategy in Stage 2 training of Self-Reflective Post-Training.
Method Stage 1 Stage 2 Val Unseen
CoT
SFT
Self-Enriched
CoT SFT
Self-Reflective
Auxiliary Task
@@
Baseline - - - 46.90 54.63 43.78
1 49.62 59.44 46.26
2 50.47 57.48 47.98
3 50.51 57.74 47.86
Full Model 51.15 59.18 48.27
Figure 3: Action decision visualization of NaviLLM [45] and our EvolveNav. We only extract two steps and display local candidate space for simplicity. Observations selected by EvolveNav (also are the ground-truth actions) and NaviLLM are annotated by green boxes and red boxes, respectively.
Effect of constructed CoT labels. Table VII compares the navigation performance under different kinds of CoT labels, where we can find the effectiveness of our introduced CoT labels by predicting landmark and direction information in a formalized way. To realize “Free-form CoT”, we introduce the free-form CoT labels collected in NavGPT-2 [43] to train the navigation agent. To realize “Direction & Landmark”, we obtain the best matched landmark in the instruction to each ground-truth observation through the CLIP model [68].
From Table VII, we can draw into the following conclusions: 1) The comparison between “Free-form CoT” and “Direction & Landmark (ours)” shows that our formalized CoT labels can effectively reduce redundant reasoning information to improve the navigational reasoning and decision accuracy. Moreover, the inference time to generate CoT at one timestep of “Free-form CoT” is $\sim$7.8s compared to $\sim$2.5s of “Direction & Landmark (ours)”, demonstrating that our method can significantly promote the reasoning speed ($\sim$$\times$3 improvement), which is crucial for sequential decision making task like navigation. 2) The comparison among “Only Direction”, “Only Landmark”, and “Direction & Landmark (ours)” reveal that both landmark and direction information are important to guide navigation decisions, demonstrating the reasonability of our constructed CoT labels. 3) The superiority of “Direction & Landmark (ours)” over “Direction & Landmark” demonstrates that our introduced CoT labels, which contain diverse landmarks, can potentially encourage the navigation agent to learn cross-modal alignment knowledge to accurately align the visual observation to the navigation instruction.
Figure 4: Visualization comparison between self-enriched chain-of-thought (CoT) labels and originally built CoT labels. Newly introduced landmarks in the self-enriched CoT label are highlighted in red fonts. GT action denotes the ground-truth action (observation). We omit the direction information in the CoT labels.
TABLE VII: Ablation study of CoT labels on Val Unseen set on R2R. denotes using the landmark mentioned in the instruction.
@@ 4.4 Visualization
In this subsection, we present various kinds of visualization results, including visualization of action decision, self-enriched CoT label, loss & performance variation, and landmark extraction, to comprehensively and deeply analyze the advantage of the proposed self-improving embodied reasoning framework.
@@
Action Decision Visualization. Fig. 3 gives the action decision visualization comparison between NaviLLM [45] and our EvolveNav, where we can find that EvolveNav generates reasonable navigational reasoning about landmarks and directions to guide correct action decision making. For example, when $t=0$, from the observations, EvolveNav infers that an observation with door and hallway represents the exit from the bedroom while NaviLLM mistakenly selects an action that remains in the bedroom. Another example is in the hallway ($t=4$ for NaviLLM and $t=1$ for EvolveNav), EvolveNav generates reasoning consistent with the decision of entering the correct doorway on the left. However, NaviLLM chooses the wrong doorway to go back. These results highlight the effectiveness of our approach in improving navigational reasoning for accurate instruction understanding and action prediction.
@@
Loss & performance variation during Stage 2 training. Fig. 5 shows the loss and performance curves during Self-Reflective Post-Training (Stage 2). In Fig. 5, we can find that our self-reflective loss $\mathcal{L}_{\mathrm{sr}}$, the total training loss $\mathcal{L}_{\mathrm{Stage2}}$, and the navigation error (NE) have similar variation trends. Especially, both two loss curves and the NE curves achieve the lowest value around epoch 9. These results show the effectiveness of our constructed self-reflective auxiliary task during Self-Reflective Post-Training in improving the navigational reasoning and decision accuracy of the LLM-based navigation agent.
Landmark extraction visualization. We compare different methods for constructing the formalized CoT labels, to verify the reasonability of our landmark extraction strategy by combining image captioning model [60] with the NLP tool Spacy [61]. Fig. 6 presents the landmark extraction visualization comparison of three image captioning models, Tag2Text [60], BLIP-v2 [69], LLaVA 1.6 vicuna 13b [70], and an open-vocabulary object recognition model RAM [71]. Concretely, we provide four candidates in a navigational step for these methods and use Spacy [61] to extract the landmarks in the output caption (except for RAM of directly obtaining tagging).
From Fig. 6, we can observe that Tag2Text can capture more informative landmarks while having less redundancy and illusion. For example, for “Cand 2”, Tag2Text correctly detects a kitchen and living room, a dining table, and chair, while BLIP-v2 only detects a large dining table. Although LLaVA 1.6 13b generates abundant captions, it brings noisy landmarks like the outdoor after noun phrases extraction. RAM also generates meaningless and non-existent landmarks, like floor and armchair. These results show that our combination of Tag2text model [60] and NLP tool for landmark extraction can effectively retain informative landmarks while reducing redundancy and illusion for constructing CoT labels, which can enable the agent to better learn cross-modal alignment between observations and instructions.
5 Conclusion
In this work, we propose EvolveNav, a novel self-improving embodied reasoning framework to fulfill generalizable and adaptable reasoning for enhancing LLM-based vision-and-language navigation. Through introducing the formalized CoT supervised fine-tuning and self-reflective post-training in the proposed framework, the agent’s navigational reasoning ability can be effectively enhanced while mitigating the overfitting to the training reasoning label distributions simultaneously to improve generalization. Experimental results on multiple VLN benchmarks under diverse training settings reveal the promising capability of our method in boosting the reasoning ability and decision accuracy for LLM-based navigation agents. We believe that our EvolveNav can provide meaningful references for designing self-improving embodied reasoning paradigms to benefit future LLM-assisted Embodied AI research.